three charges of -2.0μC, 1.0μC , and 3.0μC are in a line 12cm apart. The 3.0 μC charge is moved relative to the charge 1.0 μC until the force acting on it is zero. the final distance between 1.0μC and 3.0 μC charges is?
Answer #1

Let

Q1=-2 uC

Q2=1 uC

Q3=3 uC

Force due Q3 on Q1 =Force due to Q2 on Q3

K|Q1Q3|/0.122 =K|Q2Q1|(0.12-x)2

(0.12-X)2(3*10-6) =0.0144*(1*10-6)

0.0144-0.24x+x2=0.0048

X2-0.24x+0.0096=0

X= 0.0507 m